If the electric field in an electromagnetic wave 100 mm from a radio-emitting antenna has a maximum magnitude of 4.0 × 105 V/m what is the maximum magnetic field magnitude 600 m from the antenna?
Answer #1

Using Maxwells equation

E/B = c

B = E/c

where c is the speed of Light = 3*10^8 m/s

Electric field at 600 m = 4.0 × 10^5 V/m * 600 m = 2.4 × 10^8 V

Hence required magnetic field, B = 2.4 × 10^8 / 3*10^8 = 0.8 T